One number is 16 more than another. If the smaller number is subtracted from 2/3 of the larger, the result is 1/4 of the sum of the two numbers. find the numbers
Answer by longjonsilver(2297)

Let larger number be x
Let smaller number be y
x = y + 16 --eqn1
(2/3)x - y = (x+y)/4 --eqn2
eqn2 becomes, by multiplying every term by 12...
8x - 12y = 3x + 3y
5x = 15y
x = 3y --eqn3
sub 3 into 1 for x, to give
3y = y + 16
2y = 16
y = 8
so, x = 24
check these by using them in the question and making sure that you do get the correct answers that way.
jon